# Shadow Program Inversion with Differentiable Planning: A Framework for Unified Robot Program Parameter and Trajectory Optimization

### v1

@@ I Introduction
Intuitive robot programming has eased the use of robots to solve real-world applications. However, the cost of automation is often driven by the iterative optimization of robot trajectories and program parameters, particularly for complex manipulation tasks. This optimization is done by skilled programmers through time-consuming trial and error. “Programming by optimization” [1] allows a human programmer to specify a rough program skeleton, which is then completed by an optimization algorithm. This approach is particularly useful in tactile applications like force-controlled insertion or handling of deformable objects. However, applying general-purpose optimization methods to robot programs is challenging as the success of a robot skill depends on the parameterization of the preceding skills: An optimizer must jointly optimize the parameters of complete skill sequences or hierarchically composed subprograms. Moreover, robots must not only achieve task-level goals such as cycle time requirements, but also respect motion-level constraints such as collision-freeness or proximity to a human demonstration. Existing approaches focus exclusively on either trajectory optimization [2, 3, 4, 5] or parameter optimization [6, 7, 8, 9] and typically optimize individual skills, rather than jointly optimizing complete robot programs.
@@ III Shadow Program Inversion: A Primer
Figure 2: Left: Shadow Program Inversion with Differentiable Planning (SPI-DP) optimizes robot programs (left, red) with respect to nearly arbitrary task objectives ($\Phi$). Program parameters, such as the Force parameter of a placing action, can be optimized jointly with the low-level motion trajectories to respect task-level objectives and motion-level constraints. By performing gradient-based optimization over a differentiable surrogate (“shadow” $\bar{P}$, gray), the framework is applicable to near-arbitrary parameterized program representations, including most robot programming languages. Right: dGPMP2-ND plans collision-free motions for N-DoF serial kinematics within SPI-DP’s optimization loop. Trajectories (black) are optimized with respect to collision (blue), goal (green) and human demonstration (red) constraints, among others.
Our proposed framework is based on differentiable Shadow Program Inversion (SPI), a model-based first-order optimizer for robot program parameters [7, 39, 40], which is briefly outlined below. On the basis of SPI and a differentiable N-DoF motion planner (see Sec. IV), we present a novel double-loop first-order optimizer capable of jointly optimizing program parameters and motion trajectories (see Sec. V).
@@ III-A Differentiable Shadow Programs
The core of our framework is the concept of a shadow program: A differentiable “twin” of a robot program which serves as its surrogate for learning and optimization. The source program, which is written by the programmer and ultimately executed on the robot, and its shadow used for learning and optimization, are representationally decoupled: The source program can be expressed in any parameterized representation, such as the textual or skill-based representations used by modern robot programming frameworks [41, 42, 43].
Given a source program $P(\bm{x},\theta_{0})$, we seek to optimize the program parameters $\bm{x}$, given initial robot state $\theta_{0}$. $P$ is modeled as a function $P:\mathbb{R}^{N}\times\Theta\rightarrow\Theta^{T}$, mapping a real-valued $N$-dimensional parameter vector $\bm{x}\in\mathbb{R}^{N}$ and joint angles $\theta$ in state space $\Theta$ to the robot trajectory $\bm{\theta}\in\Theta^{T}$, where $T$ is the number of timesteps. The state space is composed of robot joint configurations and end-effector wrenches: $\Theta=\mathcal{C}\times\mathbb{R}^{6}$.
The shadow program $\bar{P}$ is a generative model of $P$, trained to approximate the real-world robot trajectory $\bm{\theta}$ for a given set of program parameters $\bm{x}$ and initial state $\theta_{0}$ [7]. A central property of $\bar{P}$ is that it is differentiable, allowing the computation of the gradient of a task-specific objective $\Phi$ over the trajectory with respect to the program’s input parameters, and, as a consequence, the optimization of $\bm{x}$ using a gradient-based optimizer.
The shadow program architecture is modular, to reflect the typically skill-based structure of most source programs. An exemplary shadow program composed of two skills is illustrated in Fig. 4. For the purpose of this paper, it is important to note that shadow skills are generative models of robot skills, predicting the expected real-world robot trajectory given the skill’s parameters and the current robot state. A differentiable planner bootstraps a naive prior trajectory, which is then refined by a neural sequence-to-sequence model to reflect expected real-world deviations from the plan (see [7] for details).
@@ III-B Robot Program Parameter Optimization
SpiralSearch and Insert are skills from a skill library, and approach_pose is a program parameter corresponding to the end-effector pose from which the robot starts searching for a hole in the workpiece. The duration of the search depends on the position of the hole relative to the approach pose. Consequently, the cycle time of this program can be optimized by adapting approach_pose to be, on average, directly above the hole. SPI solves such parameter optimization problems by first-order optimization over the shadow program, using $\bar{P}$ as a differentiable surrogate for $P$ [7]. The optimized parameters can be transferred back to the source program $P$, validated and adjusted by a human programmer, and executed on the robot.
@@ III-C Joint Parameter and Trajectory Optimization
For many tasks, program parameters and low-level motion trajectories must be jointly optimized. One example is the optimization of grasp poses to maximize grasp success: The approach and depart motions must also be optimized for collision-freeness, smoothness and other task constraints whenever the grasp pose is changed. We integrate a differentiable motion planner into the shadow program representation, which ensures that trajectories predicted in a forward pass comply with motion-level constraints such as collision-freeness, smoothness or proximity to a human demonstration. The differentiable motion planner is described in detail in Section IV.
@@ IV-A Differentiable Gaussian Motion Planning
The GP, start and goal state priors remain unchanged from the original dGPMP2 formulation, albeit extended to the N-DoF case. We contribute novel differentiable collision and joint limit factors, as well as a differentiable Cartesian demonstration prior.
@@
$(\bm{\mathcal{K}}^{-1}+\bm{H}^{T}\bm{\Sigma}^{-1}\bm{H})\delta\bm{\theta}=-\bm{\mathcal{K}}^{-1}(\bm{\theta}^{i}-\bm{\mu})-\bm{H}^{T}\bm{\Sigma}^{-1}\bm{h}(\bm{\theta}^{j})$
is solved for $\delta\bm{\theta}$, where $\bm{H}$ is the combined Jacobian, $\bm{\Sigma^{-1}}$ is the combined inverse covariance, $\bm{\mathcal{K}}^{-1}$ is the inverse kernel matrix of the GP and $\bm{h}(\bm{\theta}^{j})$ is the combined error function. All matrices are combined by concatenating the matrices of the individual factors along the row axis. The trajectory is then incrementally updated: $\bm{\theta}^{j+1}=\bm{\theta}^{j}+r*\delta\bm{\theta}$, where $r$ is the update rate.
@@ IV-C Joint Limit Factor
$h_{lim}=\begin{cases}\theta-\theta_{lim}&\ \text{if\quad}\theta>\theta_{lim}\\ -\theta_{lim}-\theta&\ \text{if\quad}\theta<-\theta_{lim}\\ 0&\ \text{otherwise}\end{cases}$
@@ IV-D Demonstration prior
Taken together, dGPMP2-ND permits collision-free motion planning by iterative optimization, while respecting additional motion-level constraints such as joint limits or adherence to a reference trajectory. dGPMP2-ND is differentiable end-to-end, permitting the differentiation of the resulting trajectories with respect to input parameters such as target poses.
@@ V Joint Trajectory and Parameter Optimization
For many real-world robot tasks, motion trajectories and program parameters cannot be optimized in isolation. Grasping is a canonical example: Grasping an object with a given grasp pose imposes constraints on the approach motion, while e.g. collision objects in the environment make some approach motions, and therefore grasp poses, impossible. Grasp poses and approach motions must be jointly optimized in order to achieve task-level objectives (a stable grasp) while obeying motion-level constraints (collision-free approach). With dGPMP2-ND, gradient-based optimization over differentiable shadow programs permits the joint optimization of motion trajectories and program parameters. Fig. 4 shows the integration of dGPMP2-ND as a differentiable collision-free motion planner into the shadow program architecture. Shadow programs are differentiable, predictive models of robot programs; with dGPMP2-ND, collision-free planning becomes part of their forward pass.
@@ V-A Shadow Program Inversion with Differentiable Planning
We find that the inner-loop dGPMP2-ND converges in less than 20 iterations for most planning problems we encountered. In each outer-loop iteration $j$, planning results are cached and used as the initial trajectories for dGPMP2-ND in the next iteration $j+1$, avoiding redundant planning. We find that one optimization iteration of a complex source program with 15 skills takes about 19 seconds on an Nvidia RTX 4090 graphics processing unit (GPU).
@@ V-B Task-Level Objective Functions
$\Phi_{cyc}(\bm{\theta})$ $=\textstyle\sum_{i=1}^{|\bm{\theta}|}\textstyle\log(\bm{\theta}_{i,EOS})$
$\Phi_{path}(\bm{\theta})$ $=\textstyle\sum_{i=2}^{|\bm{\theta}|}\lVert{\bm{\theta}}_{i,pos}-\bm{\theta}_{i-1,pos}\rVert$
$\Phi_{succ}(\bm{\theta})$ $=\textstyle\sum_{i=1}^{|\bm{\theta}|}\textstyle\log(\bm{\theta}_{i,succ})$
@@ VI-A Household Pick-and-Place with Human Demonstration
    Task-level optimization: SPI-DP can optimize the entire robot program parameters for KPIs resulting in reduced overall cycle-time while respecting imposed contact-force limits;
@@ VI-A1 Cup Pick-and-Place
A robot program consisting of approach, grasp, transfer, place and depart skills is optimized to place the cup at one of four target poses on two different shelf levels inside the cupboard. In this experiment, the approach motion is optimized to respect the motion-level constraints illustrated in Fig. 3, with a collision environment consisting of the robot, table, cup and cupboard, and a Cartesian trajectory prior given by a human demonstration. A total of 40 trials are performed, one for each combination of target pose and human demonstration. The results are shown in Fig. 5 (left). All optimized motions were collision-free, even if the human demonstration contained a collision. The target pose was reached with a mean accuracy of 0.6 mm.
@@ VI-A2 Wine Glass Pick-and-Place with Target Pose Optimization
The same robot program is optimized again, but the manipulated object is swapped for a wine glass and the gripper geometry is changed accordingly. In addition to the transfer motion, we also optimize the target pose (a parameter of the transfer skill) to minimize the cycle time of the overall program. The real-world experiment setup is shown in Fig. 5 (right). Again, 40 trials are performed, one for each combination of initial target pose and human demonstration. All optimized motions were collision-free and closely adhered to the human demonstration. SPI-DP optimized the target pose parameter resulting in motions which were 2.8 cm shorter on average. It moved the target pose as close to the shelf as possible, reducing cycle time by 40 % while avoiding collisions.
@@ VI-B Engine Block Poka-Yoke Testing with Force Control
This experiment tests the scalability of joint motion- and task-level optimization for complex industrial robot programs. The task consists of a poka-yoke quality assurance task, in which a UR5 robot arm approaches three holes on an engine block and executes a force-controlled search motion to probe the hole (see Fig. 6. To simulate stochastic process noise, the engine block is moved on a linear axis by a random offset at every iteration. SPI-DP ensures collision-free motions from one hole to another and at the same time optimizes program parameters such as target poses, search patterns, velocities and contact forces with respect to the task-level objectives of cycle time minimization and maximization of the probability of task success.
@@
20 trials are performed, for each of which the linear axis was moved randomly by up to 4 mm and the robot program was executed once with randomly initialized parameters, and once with optimized parameters. The results are shown in Table I. After optimization, the probability of finding each of the three holes is increased by 83%, 67% and 186%, respectively. In addition, optimization improved the search pattern, search dynamics and contact forces, reducing search duration by 62%. At the same time, motion-level optimization by dGPMP2-ND resulted in collision-free motions for all evaluations. The noticeably better optimization results for Hole 3 stem from its wider diameter and the greater surface area surrounding it, compared to the other two holes.
@@ VII Conclusion and Outlook
We introduce Shadow Program Inversion with Differentiable Planning (SPI-DP), a first-order robot program optimizer capable of jointly optimizing robot program parameters and motion trajectories. To that end, we present dGPMP2-ND, a differentiable motion planner for n-DoF robotic manipulators. SPI-DP integrates dGPMP2-ND into a first-order general-purpose optimizer, which optimizes the parameters of a given robot program with respect to task-level objectives, while simultaneously ensuring that motion-level constraints are respected. SPI-DP is evaluated on two representative use cases from service and industrial robotics. Both experiments show that SPI-DP enables the optimization of program parameters such as target poses or search regions while ensuring collision-freeness, smoothness and kinematic feasibility. To our knowledge, SPI-DP is the first gradient-based optimizer capable of jointly optimizing program parameters and motion trajectories for arbitrary parameterized robot programs. Limitations of SPI-DP include its relative sensitivity with respect to hyperparameters, particularly the GP and collision factor covariances. We suggest the investigation of metaheuristics [47, 48] and meta-optimization approaches [49, 50, 51] for future work, which reliably steer the optimizer toward stable solutions or optimize planner hyperparameters for efficient convergence. Moreover, we seek to evaluate dGPMP2-ND on large-scale standardized planning benchmarks and investigate its integration as a differentiable planner for reinforcement learning [27, 28, 29, 30] and Task and Motion Planning [52, 53, 22].

### v2

@@ I Introduction
Skill-based robot programming has eased the use of robots for solving real-world applications. However, the cost of automation of complex tasks is often driven by the manual tweaking of robot trajectories and program parameters, such as approach poses, waypoints or force thresholds. This labor-intensive process requires skilled programmers and time-consuming trial and error. “Programming by optimization” [1] addresses this issue by allowing a human programmer to specify a rough program skeleton, which is then completed by an optimization algorithm. This approach is particularly useful in tactile applications like force-controlled insertion or handling of deformable objects. However, applying general-purpose optimization methods to robot programs is challenging, as the success of one robot skill may depend on the parameterization of preceding skills: An optimizer must jointly optimize the parameters of complete skill sequences or hierarchically composed subprograms. Moreover, robots must not only achieve task-level goals such as cycle time requirements, but also respect motion-level constraints such as collision-freeness or proximity to a human demonstration. Existing approaches focus exclusively on either trajectory optimization [2, 3, 4, 5] or parameter optimization [6, 7, 8, 9] and typically optimize individual skills, rather than jointly optimizing complete robot programs.
@@ II-2 Trajectory optimization
Figure 2: Left: Shadow Program Inversion with Differentiable Planning (SPI-DP) optimizes skill-based robot programs (left, red) with respect to nearly arbitrary task objectives $\Phi$. Program parameters, such as the Force parameter of a placing action, can be optimized jointly with low-level motion trajectories to respect task-level objectives and motion-level constraints. By performing gradient-based optimization over a differentiable surrogate (“shadow” $\bar{P}$, gray), the framework is applicable to near-arbitrary parameterized program representations, including most robot programming languages. Right: dGPMP2-ND plans collision-free motions for N-DoF serial kinematics within SPI-DP’s optimization loop. Trajectories (black) are optimized with respect to collision (blue), goal (green), human demonstration (red) and other constraints.
@@ III Shadow Program Inversion: A Primer
Our proposed framework is based on Shadow Program Inversion (SPI), a model-based first-order optimizer for robot program parameters [7, 39, 40], which is briefly outlined below. On the basis of SPI and a differentiable N-DoF motion planner (see Sec. IV), we present a novel double-loop first-order optimizer capable of jointly optimizing program parameters and motion trajectories (see Sec. V).
@@ III-A Differentiable Shadow Programs
The core of our framework is the concept of a shadow program: A differentiable “twin” of a robot program which serves as its surrogate for learning and optimization. The source program and its shadow are representationally decoupled: The source program is written by the programmer and can be expressed in any parameterized representation, such as textual robot programming languages or task models [41, 42, 43], and is not required to be differentiable. The shadow program is a DCG, which can be automatically constructed and partially learned [7].
Given a source program $P(\bm{x},\theta_{0})$ (such as an industrial robot program in a textual programming language), parameterized by program parameters $\bm{x}$ (such as waypoints, target poses, velocities or contact forces), we seek to optimize $\bm{x}$, given the initial robot state $\theta_{0}$. Mathematically, $P$ is a function $P:\mathbb{R}^{N}\times\Theta\rightarrow\Theta^{T}$, mapping a real-valued $N$-dimensional parameter vector $\bm{x}\in\mathbb{R}^{N}$ and joint angles $\theta$ in state space $\Theta$ to the robot trajectory $\bm{\theta}\in\Theta^{T}$, where $T$ is the number of timesteps. The state space is composed of robot joint configurations and end-effector wrenches: $\Theta=\mathcal{C}\times\mathbb{R}^{6}$.
The shadow program $\bar{P}$ is a generative model of $P$, trained to approximate the real-world robot trajectory $\bm{\theta}$ for a given set of program parameters $\bm{x}$ and initial state $\theta_{0}$. The shadow program architecture is modular, to reflect the typically skill-based structure of most source programs, and permits $\bar{P}$ to be automatically constructed by traversing the source program and instantiating the corresponding shadow program skill by skill [7]. An exemplary shadow program composed of two skills is illustrated in Fig. 4. For the purpose of this paper, it is important to note that shadow skills are generative models of robot skills, predicting the expected real-world robot trajectory given the skill’s parameters and the current robot state. A differentiable planner bootstraps a naive prior trajectory, which is then refined by a neural sequence-to-sequence model to reflect expected real-world deviations from the plan (see [7] and [39] for details).
A central property of $\bar{P}$ is that it is differentiable, allowing the computation of the gradient of a task-specific objective $\Phi$ over the trajectory with respect to the program’s input parameters, and, as a consequence, the optimization of $\bm{x}$ using a gradient-based optimizer.
@@ III-B Robot Program Parameter Optimization
SpiralSearch and Insert are skills from a skill library, and approach_pose is a program parameter corresponding to the end-effector pose from which the robot starts searching for a hole in the workpiece. The cycle time of this program can be optimized by adapting approach_pose to be, on average, directly above the hole. SPI solves such parameter optimization problems by first-order optimization over the shadow program, using $\bar{P}$ as a differentiable surrogate for $P$ [7]. The optimized parameters can be transferred back to the source program $P$, validated and adjusted by a human programmer, and executed on the robot.
@@ III-C Joint Parameter and Trajectory Optimization
For many tasks, program parameters and low-level motion trajectories must be jointly optimized. One example is the optimization of grasp poses to maximize grasp success: The approach and depart motions must also be optimized for collision-freeness, smoothness and other task constraints whenever the grasp pose is changed. One of the core contributions in this paper is a differentiable motion planner, integrated into the shadow program representation, which ensures that trajectories predicted in a forward pass comply with motion-level constraints such as collision-freeness, smoothness or proximity to a human demonstration. The differentiable motion planner is described in detail in Section IV.
@@ IV-A Differentiable Gaussian Motion Planning
The GP, start and goal state priors are identical to the original dGPMP2 formulation [37], but are extended to the N-DoF case. We contribute novel differentiable collision and joint limit factors, as well as a differentiable Cartesian demonstration prior.
@@
$(\bm{\mathcal{K}}^{-1}+\bm{H}^{T}\bm{\Sigma}^{-1}\bm{H})\delta\bm{\theta}=$ $-\bm{\mathcal{K}}^{-1}(\bm{\theta}^{i}-\bm{\mu})$ (1)
$-\bm{H}^{T}\bm{\Sigma}^{-1}\bm{h}(\bm{\theta}^{j})$
is solved for $\delta\bm{\theta}$, where $\bm{H}$ is the combined Jacobian, $\bm{\Sigma^{-1}}$ is the combined inverse covariance, $\bm{\mathcal{K}}^{-1}$ is the inverse kernel matrix of the GP and $\bm{h}(\bm{\theta}^{j})$ is the combined error function [37]. All matrices are combined by concatenating the matrices of the individual factors along the row axis. The trajectory is then incrementally updated: $\bm{\theta}^{j+1}=\bm{\theta}^{j}+r*\delta\bm{\theta}$, where $r$ is the update rate.
@@ IV-C Joint Limit Factor
$h_{lim}=\begin{cases}\theta-\theta_{lim}&\ \text{if\quad}\theta>\theta_{lim}\\ -\theta_{lim}-\theta&\ \text{if\quad}\theta<-\theta_{lim}\\ 0&\ \text{otherwise}\end{cases}$ (2)
@@ IV-D Demonstration Prior
Taken together, dGPMP2-ND permits collision-free motion planning by iterative optimization, while respecting additional motion-level constraints such as joint limits or adherence to a reference trajectory. dGPMP2-ND permits the differentiation of planned trajectories with respect to input parameters such as target poses, as well as the integration into second- or higher-order optimizers. In our implementation, gradients are computed via automatic differentiation [45].
@@ V Joint Trajectory and Parameter Optimization
For many real-world robot tasks, motion trajectories and program parameters cannot be optimized in isolation. Grasping is a canonical example: Grasping an object with a given grasp pose imposes constraints on the approach motion, while e.g. collision objects in the environment make some approach motions, and therefore grasp poses, impossible. Grasp poses and approach motions must be jointly optimized in order to achieve task-level objectives (a stable grasp) while obeying motion-level constraints (collision-free approach). With dGPMP2-ND, gradient-based optimization over differentiable shadow programs permits the joint optimization of motion trajectories and program parameters. Fig. 4 shows the integration of dGPMP2-ND as a differentiable collision-free motion planner into the shadow program architecture. Shadow programs are differentiable, predictive models of robot programs; with dGPMP2-ND, collision-free planning is part of their forward pass.
@@ V-A Shadow Program Inversion with Differentiable Planning
Note that not all shadow skills require a computationally demanding planner such as dGPMP2-ND for generating prior trajectories. For purely Cartesian skills such as spiral search [39] or linear motions, simple Cartesian planners suffice [7]. We find that the inner-loop dGPMP2-ND converges in less than 20 iterations for most planning problems we encountered. In each outer-loop iteration $j$, planning results are cached and used as the initial trajectories for dGPMP2-ND in the next iteration $j+1$, avoiding redundant planning. We find that one optimization iteration of a complex source program with 15 skills takes about 19 seconds on an Nvidia RTX 4090 graphics processing unit (GPU).
@@ V-B Task-Level Objective Functions
$\Phi_{cyc}(\bm{\theta})$ $=\textstyle\sum_{i=1}^{|\bm{\theta}|}\textstyle\log(\bm{\theta}_{i,EOS})$ (3)
$\Phi_{path}(\bm{\theta})$ $=\textstyle\sum_{i=2}^{|\bm{\theta}|}\lVert{\bm{\theta}}_{i,pos}-\bm{\theta}_{i-1,pos}\rVert$ (4)
$\Phi_{succ}(\bm{\theta})$ $=\textstyle\sum_{i=1}^{|\bm{\theta}|}\textstyle\log(\bm{\theta}_{i,succ})$ (5)
@@ VI-A Household Pick-and-Place with Human Demonstration
    Task-level optimization: SPI-DP can optimize the entire robot program parameters with respect to task-level objectives, resulting in reduced overall cycle-time while respecting imposed contact-force limits;
@@ VI-A1 Cup Pick-and-Place
A robot program consisting of approach, grasp, transfer, place and depart skills is optimized to place the cup at one of four target poses on two different shelves inside the cupboard. In this experiment, the approach motion is optimized to respect the motion-level constraints illustrated in Fig. 3, with a collision environment consisting of the robot, table, cup and cupboard, and a Cartesian trajectory prior given by a human demonstration. A total of 40 trials are performed, one for each combination of target pose and human demonstration. The results are shown in Fig. 5 (left). All optimized motions were collision-free, even if the human demonstration contained a collision. The target pose was reached with a mean accuracy of 0.6 mm. The demonstration primarily acts as a regularizer that enforces implicit motion constraints, such as keeping the cup upright.
@@ VI-A2 Wine Glass Pick-and-Place with Target Pose Optimization
The same robot program is optimized again, but the manipulated object is swapped for a wine glass and the gripper geometry is changed accordingly. In addition to the transfer motion, we also optimize the target pose (a parameter of the transfer skill) to minimize the cycle time of the overall program. The real-world experiment setup is shown in Fig. 5 (right). Again, 40 trials are performed, one for each combination of initial target pose and human demonstration. All optimized motions were collision-free. SPI-DP optimized the target pose parameter of the transfer skill to be as close to the shelf as possible, yielding placing motions that were 2.8 cm shorter on average than before optimization. The placing motion has outsize impact on the cycle time due to its slow speed, which is required for force-controlled placing; overall, average cycle times were reduced by 40 % while avoiding collisions.
@@ VI-B Engine Block Poka-Yoke Testing with Force Control
This experiment tests the scalability of joint motion- and task-level optimization for complex industrial robot programs. The task consists of a poka-yoke quality assurance task, in which a UR5 robot arm approaches three holes on an engine block and executes a force-controlled spiral search motion to probe the hole (see Fig. 6. To simulate stochastic process noise, the engine block is moved on a linear axis by a random offset at every iteration. SPI-DP ensures collision-free motions from one hole to another and at the same time optimizes program parameters such as target poses, search patterns, velocities and contact forces with respect to the task-level objectives of cycle time minimization and maximization of the probability of task success (dropping into the hole before exausting the search pattern).
@@
20 trials are performed, for each of which the linear axis was moved randomly by up to 4 mm and the robot program was executed once with randomly initialized parameters, and once with optimized parameters. After optimization, the probability of finding each of the three holes is increased by 83%, 67% and 186%, respectively (see Table I). In addition, optimization improved the search pattern, search dynamics and contact forces, reducing search duration by 62%. Motions remained collision-free throughout. The noticeably better optimization results for hole 3 stem from its wider diameter and the greater surface area surrounding it, compared to the other holes.
@@ VII Conclusion and Outlook
We introduce Shadow Program Inversion with Differentiable Planning (SPI-DP), a first-order robot program optimizer capable of jointly optimizing program parameters and motion trajectories. We present dGPMP2-ND, a differentiable motion planner for N-DoF serial manipulators. SPI-DP leverages dGPMP2-ND to optimize the parameters of robot programs with respect to task-level objectives, while simultaneously enforcing motion-level constraints. Experiments on two representative use cases from service and industrial robotics show that SPI-DP optimizes program parameters such as target poses or search regions while ensuring collision-freeness, smoothness and kinematic feasibility. To our knowledge, SPI-DP is the first gradient-based optimizer capable of jointly optimizing program parameters and motion trajectories for arbitrary parameterized robot programs. Limitations include its relative sensitivity to hyperparameters, particularly the GP and collision factor covariances. We suggest the future investigation of metaheuristics [47, 48] and meta-optimization [49, 50, 51] to steer the optimizer toward stable solutions or optimize hyperparameters for efficient convergence. Moreover, we seek to evaluate dGPMP2-ND on large-scale planning benchmarks and investigate its integration as a differentiable planner for reinforcement learning [27, 28, 29, 30] and Task and Motion Planning [52, 53, 22].